Problem 11-28 Determining the break-even point and preparing a contribution margin income statement LO 11-5 Ritchie Manufacturing Company makes a product that it sells for $150 per unit. The company incurs variable manufacturing costs of $64 per unit. Variable selling expenses are $17 per unit, annual fixed manufacturing costs are $494,000, and fixed selling andd administrative costs are $237,400 per year. Required Determine the break-even point in units and dollars using each of the following approaches: a. Use the equation method b. Use the contribution margin per unit approach. c. Prepare a contribution margin income statement for the break-even sales volume. Complete this question by entering your answers in the tabs below. Req A to B | Req C Determine the break-even point in units and dollars using the equation method, the contribution margin per unit approach and the contribution margin ratio approach. a. Break-even point in units Break-even point in dollars b. Contribution margin per unit Break-even point in units Break-even point in dollars Req A to B Req CProblem 11-28 Determining the break-even point and preparing a contribution margin income statement LO 11-5 Ritchie Manufacturing Company makes a product that it sells for $150 per unit. The company incurs variable manufacturing costs of $64 per unit. Variable selling expenses are $17 per unit, annual fixed manufacturing costs are $494,000, and fixed selling and administrative costs are $237,400 per year Required Determine the break-even point in units and dollars using each of the following approaches: a. Use the equation method b. Use the contribution margin per unit approach. c. Prepare a contribution margin income statement for the break-even sales volume. Complete this question by entering your answers in the tabs below. Req A to B Req C Prepare a contribution margin income statement for the break-even sales volume RITCHIE MANUFACTURING COMPANY Contribution Margin Income Statement
Answer #1

a. Use the equation method

Assume the break even units to be "x".

We know that at break even level of sales, the profit will be zero -

Sales - Variable Costs - Fixed Costs = 0

150x - (64 + 17)x - (494,000 + 237,400) = 0

150x - 81x = $731,400

x = 10,600 units

Break even point in units = 10,600 units

Break even point in dollars = 10,600 units * $150 = $1,590,000

b. Use the contribution margin per unit approach

Contribution margin per unit = Selling price per unit (-) Variable cost per unit = $150 - ($64 + $17) = $69 per unit

Break even point in units = Fixed costs / Contribution margin per unit = (494,000 + 237,400) / 69 = 10,600 units

Break even point in dollars = 10,600 units * $150 = $1,590,000

c. Contribution Margin Income Statement for break even sales volume

Income Statement at break even sales volume of 10,600 units -

Calculation $1,590,000 10,600 $150 Selling price Variable cost uring (678,400) 10,600 $64 Manufacturing Selling( Contribution p.u Fixed cost ($180,200) 10,600* $17 $731,400 Manufacturing Selling ($494,000) ($237,400) Profit $0